Joe is 3 times as old as Mack. The difference in their age is 24
Found 2 solutions by Maths68, oberobic:
Answer by Maths68(1474)

Let
Present age of Joe = j
Present age of Mack = m
Given
Joe is 3 times as old as Mack
j=3m ........(1)
The difference in their age is 24
j-m=24
Put the value of j in above equation we have
3m-m=24
2m=24
Divide by m both sides of the equation
2m/2=24/2
m=12
Put the value of m in (1)
j=3m
j=3(12)
j=36
Present age of Joe = 36 years old
Present age of Mack = 12 years old

Answer by oberobic(2304)

J = Joe's age
M = Mack's age
.
J = 3*M
.
J -M = 24
so
J = 24 + M
.
substitute J = 24+M
.
24+M = 3M
2M = 24
M = 12
.
J =3(12)
J =36
.
The difference in their ages is 36-12=24.
.
Answer: Joe is 36, and Mack is 12.
.
Done.